OpenStudy (anonymous):

Find the equation of the tangent line for the given function at x=(pi/3). f(x)=4tan(x)

OpenStudy (anonymous):

So I already worked this problem out, and found that the slope of the tangent line is 16.

OpenStudy (anonymous):

I give the equation of the tangent line using y-y1=m(x-x1)

OpenStudy (anonymous):

But I think I'm missing something. Is there a y1 value here? I don't see one.

OpenStudy (anonymous):

x1=(pi/3), m=(16), y1=(?)

OpenStudy (anonymous):

I thought that y would be zero, but wolfram is telling me that is incorrect.

OpenStudy (anonymous):

y = 4tan(pi/3)

OpenStudy (anonymous):

Oh...psshh. THat makes sense lol.

OpenStudy (anonymous):

I get stupid sometimes.

OpenStudy (anonymous):

we all do. http://openstudy.com/updates/506e1f6ee4b060a360ff3a55

OpenStudy (anonymous):

algebra fail lol.

OpenStudy (anonymous):

Ha, yep. Thanks.